# AstroMLab 4: Benchmark-Topping Performance in Astronomy Q&A with a 70B-Parameter Domain-Specialized Reasoning Model

### v1

@@ IV Evaluation
Figure 3: Performance comparison on the AstroMLab-1 benchmark across 38 LLMs as of May 2025. The x-axis shows cost per 0.1M tokens (USD) on a logarithmic scale, while the y-axis shows accuracy percentage. AstroSage-Llama-3.1-70B achieves 86.2%, outperforming all other models including more expensive proprietary offerings like o3, Claude-3.7-Sonnet, and GPT-4.1. The diagonal dashed lines represent cost-efficiency trade-offs, where a tenfold increase in cost typically corresponds to a 3.5 percentage point improvement in accuracy. The vertical red arrows highlight the effect of domain specialization, showing that both AstroSage models jump approximately two cost-efficiency lines compared to their base models (Llama-3.1-8B and Llama-3.1-70B), representing a roughly 100×100$\times$100 × improvement in cost-efficiency. The Wilson Score interval in the bottom right shows the typical uncertainty due to the finite number of questions. Figure adapted from Ting et al. [1].
To evaluate the performance of AstroSage-Llama-3.1-70B, we utilize the AstroMLab-1 benchmark developed in Ting et al. [1]. This benchmark consists of 4,425 high-quality, human-verified multiple-choice questions spanning astronomy, astrophysics, cosmology, and astronomical instrumentation. Importantly, these questions are derived from Annual Review of Astronomy and Astrophysics (ARAA) papers that were explicitly withheld from the AstroSage training corpus. This ensures the model is evaluated on genuinely unseen material, and its performance is not merely an artifact of training on the benchmark’s source texts.
On this benchmark, AstroSage-Llama-3.1-70B achieves a score of $86.2\%$ without enabling reasoning. As illustrated in Figure 3, this performance establishes AstroSage-Llama-3.1-70B as the leading model, outperforming all other tested open-weight and proprietary models. This includes a substantial improvement over its 8-billion parameter predecessor, AstroSage-Llama-3.1-8B, as well as contemporary large-scale general-purpose LLMs such as o3, GPT-4.1, Claude-3.7-Sonnet, and Deepseek-R1. For context, as mentioned in TdH25, professional astronomers score around 67% on this benchmark.
Our evaluation substantially updates the results originally presented in Ting et al. [1], which was published in July 2024. The benchmark analysis includes a cost-accuracy trade-off visualization, represented by diagonal dashed lines in Figure 3. This analysis reveals that within a model family (e.g., GPT-4.1 series or Gemini series), a tenfold increase in API cost typically corresponds to an improvement of approximately 3.5 percentage points in accuracy. Consequently, the distance between adjacent lines represents an order of magnitude gain in cost efficiency.
As highlighted by the vertical red arrows in Figure 3, our domain specialization approach delivers remarkable efficiency gains. Both AstroSage models jump approximately two cost-efficiency lines compared to their respective base models (Llama-3.1-8B and Llama-3.1-70B), representing improvement of approximately 100×100$\times$100 × in cost-efficiency. This demonstrates that targeted domain specialization can achieve performance levels that would otherwise require models costing two orders of magnitude more at inference time.
For testing, we provided all models with minimal system prompts indicating they should act as astronomy experts. This approach ensures homogeneous evaluation conditions across all models. We verified that for the high-performing models included in our comparison, variables such as temperature settings played a relatively minor role in performance outcomes, further confirming the robustness of our comparative analysis.
The original study predicted model improvements would shift performance to the next diagonal line every three to six months, a forecast that has proven accurate. Most leading models from July 2024 performed at the level where Step-2111The Step-2 results presented here reflect the version of the model as of November 2024, which differs from the version evaluated in the original Ting et al. [1] study in July 2024. now resides, and we observe that current models have largely surpassed this efficiency level. A more detailed analysis of these trends will be presented in a forthcoming paper dedicated to benchmarking efforts.
An interesting observation from our updated evaluation is that while the 3.5 percentage point trade-off slope still holds for some series (notably the Gemini-2.0 and Gemini-2.5 series), models like Qwen-3 and GPT-4.1 exhibit steeper drop-offs in performance across their tiered offerings. This suggests that current distillation approaches for creating more affordable variants of powerful models may be less effective for specialized knowledge domains like astronomy.
For our cost analysis, we took the average of the cost of input tokens and the cost of output tokens, as listed at https://artificialanalysis.ai/leaderboards/providers at the time of testing. While these costs may fluctuate by tens of percentage points, such variations would not significantly alter the comparative landscape presented in Figure 3.
In our evaluation methodology, we applied a consistent approach to models with reasoning capabilities. For models with explicit reasoning modes (e.g., Gemini-2.5-Flash), we enabled this feature during testing. While this approach may not yield perfectly equivalent cost comparisons due to the increased token usage in reasoning models, it provides a standardized basis for performance assessment. Interestingly, we found that enabling reasoning modes generally did not significantly improve scores for most models on this benchmark, including AstroSage-Llama-3.1-70B. This finding may be due to the AstroMLab-1 questions primarily testing System 1 thinking (fast, intuitive knowledge recall) rather than the complex multi-step thought process where reasoning modes typically demonstrate advantages.
@@
We acknowledge this as a limitation of current astronomy benchmarks, which offer limited evaluation of problem-solving capabilities requiring deep reasoning. Nevertheless, since a primary goal of our specialized training is to imbue the model with comprehensive domain knowledge, we consider this benchmark a reasonable test of our model’s capabilities. Future work should develop more complex reasoning-based astronomy benchmarks to better assess the full potential of reasoning-enhanced models in this domain.
The exceptional performance of AstroSage-Llama-3.1-70B on this benchmark supports our central hypothesis: domain specialization, when scaled to larger models, can enable specialized models to surpass even the most advanced general-purpose commercial models within their domain of expertise.
V Availability
@@
To encourage adoption, further research, and community-driven improvements, AstroSage-Llama-3.1-70B is made freely available under the Llama 3.1 Community License. The full model weights can be accessed and downloaded from our project repository on Hugging Face: https://huggingface.co/AstroMLab/AstroSage-70B.
@@ VI Discussion and Conclusion
The development of AstroSage-Llama-3.1-70B advances specialized language models for astronomy. Building on the foundation established by AstroSage-Llama-3.1-8B, this 70-billion parameter model incorporates a more powerful base architecture, enriched training datasets, refined training methodologies, and explicit reasoning capabilities. In the Ting et al. [1] benchmark, AstroSage-Llama-3.1-70B achieved state-of-the-art performance with a score of 86.2%, outperforming all other models at the time of writing, including proprietary models that are orders of magnitude more expensive to operate.
@@
Our results support the central hypothesis of this work: domain specialization, when scaled to larger models, can enable specialized systems to surpass even the most advanced general-purpose commercial models within their domain of expertise. However, it is crucial to note that this superior performance on astronomy-specific tasks does not necessarily translate to general capabilities. In tasks requiring complex mathematical reasoning or general problem-solving, more generalized models likely maintain substantial advantages. Nevertheless, through internal testing and preliminary expert evaluations, we have observed that AstroSage-Llama-3.1-70B demonstrates enhanced domain knowledge recall and the ability to make connections between astronomical concepts compared to general-purpose models. While these qualitative assessments are promising, we acknowledge that more rigorous, large-scale arena-type testing would be necessary to fully validate these observations.
An interesting trend emerged in our evaluation regarding the effectiveness of model distillation and scaling. The performance drop-off between flagship models and their smaller variants appears more pronounced than observed a year ago, particularly in the 30-70B parameter range. This is especially evident in models like Qwen-3, where performance decreases more steeply than the historical 3.5 percentage point per order of magnitude cost relationship would predict. This trend becomes even more pronounced at the 8B parameter level, highlighting the potential importance of specialized training for deploying cost-effective models at scale.
@@
The practical implications of these findings for astronomy are still emerging. While AstroSage-Llama-3.1-70B at 70B parameters represents a substantial computational investment, it may prove valuable for tasks that require both scalability and domain expertise, such as processing and summarizing large volumes of astronomical literature. The open-weight nature of AstroSage-Llama-3.1-70B also enables offline processing of proprietary data from institutions like NASA and NSF, addressing a practical consideration for many research applications.
Looking forward, two key areas warrant further investigation:
1. Development of more comprehensive benchmarks that specifically evaluate reasoning capabilities in astronomical contexts, including problem-solving tasks that more closely resemble real research challenges.
2. Integration of AstroSage-Llama-3.1-70B with astronomy-specific tools and workflows, moving toward more comprehensive AI research assistants that can handle both domain knowledge and practical research tasks.
@@
AstroSage-Llama-3.1-70B advances the integration of AI assistants into astronomical research and education. By making such powerful, specialized tools openly available, we aim to democratize access to cutting-edge AI and accelerate scientific discovery. As astronomical data continues to grow in volume and complexity, tools like AstroSage-Llama-3.1-70B will become increasingly valuable for extracting insights, generating hypotheses, and accelerating the pace of scientific discovery.
@@ Acknowledgements
This research used resources of the Oak Ridge Leadership Computing Facility (OLCF), which is a DOE Office of Science User Facility at the Oak Ridge National Laboratory supported by the U.S. Department of Energy under Contract No. DE-AC05-00OR22725 and support from Microsoft’s Accelerating Foundation Models Research (AFMR) program. TdH was supported by World Premier International Research Center Initiative (WPI), MEXT, Japan. YST is supported by the National Science Foundation under Grant No. 2406729. Work at Argonne National Laboratory is supported by UChicago Argonne LLC, Operator of Argonne National Laboratory. Argonne, a U.S. Department of Energy Office of Science Laboratory, is operated under Contract No. DE-AC02-06CH11357. A special thanks to Sajal Dash for suggesting the use of the GPT-NeoX LLM Training framework.

### v2

@@ 4.2 Model Performance
Figure 3: Performance comparison on the AstroMLab-1 benchmark. The x-axis shows cost per 0.1M tokens (USD) on a logarithmic scale; the y-axis shows accuracy percentage. Open-weight models are shown as triangles; proprietary models as circles. The diagonal dashed lines represent cost-efficiency trade-offs where a tenfold increase in cost corresponds to approximately 3.5 percentage points improvement. The Wilson Score interval shows uncertainty due to finite sample size of questions.
On this benchmark, AstroSage-Llama-3.1-70B achieves a score of $89.0\%$, placing it among the top models alongside GPT-5.2 (89.4%), Claude 4.5 Opus (89.0%), and Gemini 3.0 Pro (88.9%)—all released in late 2025, months after AstroSage-Llama-3.1-70B was finalized in May 2025. Gemini 3.0 Flash (90.3%) is the only model to surpass AstroSage-Llama-3.1-70B in absolute score; despite being a “Flash” variant, it outperforms even its own Pro version. However, Gemini 3.0 Flash achieves a lower value score (+4.98 vs. +5.55) due to higher inference cost—in practical terms, processing the entire astro-ph archive would cost approximately \$10,000 with Gemini 3.0 Flash versus \$1,000 with AstroSage-Llama-3.1-70B. For context, professional astronomers score around 67% on this benchmark (de_haan_achieving_2025).
Figure 3 illustrates the cost-performance landscape. For cost analysis, we averaged input and output token costs as listed at https://llmtokencost.com, cross-verified with https://artificialanalysis.ai. For AstroSage models, we assume the same inference cost as Llama-3.1 models of equivalent size, as they share identical architectures. While API pricing changes over time, such variations do not qualitatively alter the comparative landscape. As identified in tingAstroMLab1Who2024, leading models cluster along diagonal lines where a tenfold cost increase corresponds to roughly 3.5 percentage points in accuracy—a relationship that persists even as the frontier has shifted upward since summer 2024.
4.3 Temporal Evolution
[…]
The gap between open-weight and proprietary frontiers has narrowed. While proprietary models led through early 2024, open alternatives such as DeepSeek-R1, Qwen-3, and the GPT-OSS series now approach parity. This is relevant for astronomy, where open models enable deployment on institutional infrastructure and processing of sensitive data.
[…]
An emerging trend is that performance gaps between flagship models and their smaller variants have widened. For example, Claude 4.5 Opus (89.0%) outperforms Claude 4.5 Haiku (83.3%) by 5.7 percentage points, and GPT-5.2 (89.4%) exceeds GPT-5-nano (83.1%) by 6.3 points. This widening gap reinforces the value of domain-specific training for deploying cost-effective models.
@@ 4.3 Temporal Evolution
We acknowledge that current astronomy benchmarks offer limited evaluation of problem-solving capabilities requiring deep reasoning. Future work should develop more complex reasoning-based astronomy benchmarks to better assess the full potential of reasoning-enhanced models in this domain.
5 Availability
@@
To encourage adoption, further research, and community-driven improvements, AstroSage-Llama-3.1-70B is made freely available under the Llama 3.1 Community License. The full model weights can be accessed and downloaded from our project repository on Hugging Face: https://huggingface.co/AstroMLab/AstroSage-70B-20251009.
@@ 6 Discussion and Conclusion
The development of AstroSage-Llama-3.1-70B advances specialized language models for astronomy. Building on the foundation established by AstroSage-Llama-3.1-8B, this 70-billion parameter model incorporates a more powerful base architecture, enriched training datasets, refined training methodologies, and explicit reasoning capabilities. On the AstroMLab-1 benchmark, AstroSage-Llama-3.1-70B achieved a score of $89.0\%$, placing it among the top-performing models alongside advanced proprietary offerings, while being more cost-efficient.
@@
Our results support the central hypothesis of this work: domain specialization, when scaled to larger models, can enable specialized systems to compete with advanced general-purpose commercial models within their domain of expertise. However, it is crucial to note that this strong performance on astronomy-specific tasks does not necessarily translate to general capabilities. In tasks requiring complex mathematical reasoning or general problem-solving, more generalized models likely maintain advantages. Nevertheless, through internal testing and preliminary expert evaluations, we have observed that AstroSage-Llama-3.1-70B demonstrates improved domain knowledge recall and the ability to make connections between astronomical concepts compared to general-purpose models. While these qualitative assessments are promising, we acknowledge that more rigorous, large-scale arena-type testing would be necessary to fully validate these observations.
@@
The practical implications of these findings for astronomy are still emerging. While AstroSage-Llama-3.1-70B at 70B parameters represents a computational investment, it may prove valuable for tasks that require both scalability and domain expertise, such as processing and summarizing large volumes of astronomical literature. The open-weight nature of AstroSage-Llama-3.1-70B also enables offline processing of proprietary data from institutions like NASA and NSF, addressing a practical consideration for many research applications.
Future development must address two emerging frontiers. First, the field requires benchmarks that transcend simple knowledge retrieval. To fully assess models like AstroSage-Llama-3.1-70B, we must develop evaluation suites that test multi-step reasoning, hypothesis formulation, and data interpretation in realistic research scenarios. Second, the potential of this model will be best realized through integration with astronomical software pipelines. Moving beyond a chat interface to an agentic framework—where the model can query databases (e.g., SIMBAD, ADS) or execute analysis code—represents the next leap in AI-assisted discovery.
@@
AstroSage-Llama-3.1-70B advances the integration of AI assistants into astronomical research and education. By making such specialized tools openly available, we aim to democratize access to cutting-edge AI and accelerate scientific discovery. As astronomical data continues to grow in volume and complexity, tools like AstroSage-Llama-3.1-70B will be valuable for extracting insights, generating hypotheses, and accelerating the pace of scientific discovery.
@@ Acknowledgements
This research used resources of the Oak Ridge Leadership Computing Facility (OLCF), which is a DOE Office of Science User Facility at the Oak Ridge National Laboratory supported by the U.S. Department of Energy under Contract No. DE-AC05-00OR22725 and support from Microsoft’s Accelerating Foundation Models Research (AFMR) program. TdH was supported by World Premier International Research Center Initiative (WPI), MEXT, Japan. YST is supported by the National Science Foundation under Grant No. 2406729 and a Humboldt Research Award from the Alexander von Humboldt Foundation. Work at Argonne National Laboratory is supported by UChicago Argonne LLC, Operator of Argonne National Laboratory. Argonne, a U.S. Department of Energy Office of Science Laboratory, is operated under Contract No. DE-AC02-06CH11357. The ongoing public deployment of AstroSage-Llama-3.1-70B is supported by ACCESS allocation PHY240259. A special thanks to Sajal Dash for suggesting the use of the GPT-NeoX LLM Training framework.